Find the x-intercept and y-intercept of 2x+3y=-18. Thankyou
Found 2 solutions by nyc_function, CharlesG2:
Answer by nyc_function(2741)

HINT:
To find the x-intercept, replace y with 0 and solve for x.
To find the y-intercept, replace x with 0 and solve for y.
Can you do it now?

Answer by CharlesG2(834)

Find the x-intercept and y-intercept of 2x+3y=-18. Thankyou
x-intercept or horizontal intercept is when the line crosses the x-axis,
in other words when y=0
y-intercept or vertical intercept is when the line crosses the y-axis,
in other words when x=0 or the point (0,b) where b is the y-intercept
2x + 3y = -18
(put in form y=mx + b, where m is the slope, and b is the y-intercept)
3y = -2x - 18
y = (-2/3)x - 6
your slope m is -2/3
your y-intercept b is 6 or (0,6)
plug in y=0 for y and solve for x to get x-intercept
0 = (-2/3)x - 6
6 = (-2/3)x
6 * (-3/2) = x
-18/2 = -9 = x
your x-intercept is -9 or (-9,0)